### v1

# Deep Unlearn: Benchmarking Machine Unlearning

Abstract
Machine unlearning (MU) aims to remove the influence of particular data points from the learnable parameters of a trained machine learning model. This is a crucial capability in light of data privacy requirements, trustworthiness, and safety in deployed models. MU is particularly challenging for deep neural networks (DNNs), such as convolutional nets or vision transformers, as such DNNs tend to memorize a notable portion of their training dataset. Nevertheless, the community lacks a rigorous and multifaceted study that looks into the success of MU methods for DNNs. In this paper, we investigate 18 state-of-the-art MU methods across various benchmark datasets and models, with each evaluation conducted over 10 different initializations, a comprehensive evaluation involving MU over 100K models. We show that, with the proper hyperparameters, Masked Small Gradients (MSG) and Convolution Transpose (CT), consistently perform better in terms of model accuracy and run-time efficiency across different models, datasets, and initializations, assessed by population-based membership inference attacks (MIA) and per-sample unlearning likelihood ratio attacks (U-LiRA). Furthermore, our benchmark highlights the fact that comparing a MU method only with commonly used baselines, such as Gradient Ascent (GA) or Successive Random Relabeling (SRL), is inadequate, and we need better baselines like Negative Gradient Plus (NG+) with proper hyperparameter selection.
1 Introduction
Machine unlearning aims to remove the influence of a specified subset of training data points from trained models [10]. This process is crucial for enhancing privacy preservation, model safety, and overall model quality. MU helps ensure compliance with the right to be forgotten [23], removes erroneous data points that negatively impact model performance [43], and eliminates biases introduced by parts of the training data [14]. Deep neural networks present significant challenges for MU due to their computationally intensive training requirements, highly non-convex loss landscapes, and tendency to memorize substantial portions of their training data [22, 12, 21]. The key open challenges in MU include: (1) A degradation in model accuracy often accompanies unlearning; (2) Some MU methods require the model to be trained in specific ways, such as saving checkpoints, tracking accumulated gradients, or training with differential privacy, limiting their applicability to already deployed models; (3) Assurance of information removal is difficult as there are no reliable metrics to measure it accurately; (4) There is no consensus on which methods are the most effective.
A branch of MU known as exact unlearning aims to guarantee that the specified forget data have been completely removed from the model. The most reliable exact MU method is to retrain the model from scratch while excluding the forget data. Another exact MU that offers data removal guarantees is SISA (Sharded, Isolated, Sliced, Aggregated) [9]. However, exact MU is computationally prohibitive, emphasizing the need for more efficient methods. The alternative branch is approximate unlearning that aims to approximate data deletion and is often less precise but more computationally efficient. Approximate MU lacks theoretical guarantees, necessitating empirical evaluations to determine their effectiveness, reliability, and computational efficiency across various datasets. Thus, the community needs a proper benchmarking and evaluation of state-of-the-art approximate MU methods.
[…]
The contributions of our study is to address the following research questions:
Q1. Are the commonly used MU baselines reliable? Most recently introduced MU methods are compared only with three baselines: Fine-tuning, Gradient Ascent, and Successive Random Labels, and not across recently introduced approaches. We show that with proper hyperparameter selection FT is a reliable baseline. In contrast, GA consistently performs poorly and should be replaced by the more recent Negative Gradient Plus (NG+) [37] that simultaneously reduces the performance in the forget set while maintaining the performance on the rest of the data points.
Q2. How reliable are MU methods across datasets, models, and initializations? Our findings show that, unlike most recent MU methods, Masked-Small-Gradients is consistently among the best performing methods across various metrics. In contrast, methods such as SRL do not consistently outperform others across different datasets.
[…]
2 Background of Machine Unlearning
Setting. Starting with a training set $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ and a trained model $f_{O}$, referred to as the original model, the objective of a MU method $U$ is to remove the influence of a particular subset of training set $\mathcal{D}_{F}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{K}\subset\mathcal{D}$ referred to as the forget set where $K\ll N$. The rest of training set $\mathcal{D}_{R}=\mathcal{D}\backslash\mathcal{D}_{F}$ is called the retain set. The forget set and the retain set are distinct and complementary subsets of the training set. The outcome of MU is an unlearned model $f_{U}$, the aim for which is to perform on par with a model retrained from scratch on $\mathcal{D}_{R}$; this latter model $f_{R}$ is referred to as the retrained model. We denote the weights of the original model and the retrained model as $\theta_{O}$ and $\theta_{R}$, respectively. For evaluations, we consider two held-out sets: the validation set $\mathcal{D}_{V}$ and test set $\mathcal{D}_{T}$, both drawn from the same distribution as $\mathcal{D}$. We consider the accuracy of the retrained model as the optimal accuracy. One critical assumption we make is that the MU method has access to $\theta_{O}$, $\mathcal{D}_{R}$, $\mathcal{D}_{F}$, and $\mathcal{D}_{V}$.
(1) Unlearning Evaluation. To evaluate the success of unlearning, one approach is to check whether data points in $\mathcal{D}_{F}$ still influence the predictions made by the unlearned model  [13, 37, 31]. This is commonly done via influence functions [35, 8, 29], membership inference attacks (MIA) [44]. MIA has become one of the most common approaches for evaluating MU methods. It aims to determine whether specific data points were part of the original training dataset based on the unlearned model. We compute the population-based U-MIA, denoted with $\text{MIA}(\mathcal{D},f_{U})$ evaluated on data $\mathcal{D}$ [^1] The methodology for this follows the classic MIA: we compute the losses on $\mathcal{D}_{F}$ and $\mathcal{D}_{V}$, we shuffle and trim them so that they are of equal size. We then train logistic regression models in a 10-fold cross validation, and compute the average accuracy across the folds.. Using this we define the discernibility metric as $\text{Disc}(\mathcal{D}_{V},f_{U})=|2\times\text{MIA}(\mathcal{D},f_{U})-1|\in[0,1]$, and similarly, indiscernibility is given by $\text{Indisc}(\mathcal{D},f_{U})=1-\text{Disc}(\mathcal{D},f_{U})\in[0,1]$. We set $\mathcal{D}$ to either the test set $\mathcal{D}_{T}$ or validation set $\mathcal{D}_{V}$. The indiscernibility equals $1$ when the accuracy of the MIA is not better than random guessing. A more recent MIA variation is the unlearning likelihood ratio attack (U-LiRA) [37]. As highlighted by [31], U-LiRA is a more robust evaluation approach for approximate MU. Nonetheless, U-LiRA is much more computationally demanding than U-MIA. We evaluate the methods that defeat the weaker, less expensive U-MIA attack, additionally against the U-LiRA attack.
(2) Accuracy. The classification accuracy of unlearned model on the retain set should be as close as possible to that of the original model. First, we consider three metrics derived directly from the model’s classification accuracy on different sets: retain accuracy (RA), forget accuracy (FA), and test accuracy (TA). RA is defined as
$\text{RA}(\mathcal{D}_{R},f_{U})=\frac{1}{|\mathcal{D}_{R}|}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{D}_{R}}\mathbf{1}_{y_{i}=f_{U}(\mathbf{x}_{i})}\in[0,1].$ (1)
The metrics for FA and TA can be derived by replacing $\mathcal{D}_{R}$ with $\mathcal{D}_{F}$ and $\mathcal{D}_{T}$, respectively. Second, while RA, FA, and TA give us insight into the overall accuracy of the unlearned model, they do not capture how well it performs compared to a retrained model $f_{R}$. Considering the $f_{R}$ model as a gold standard, we derive three more metrics: retain retention (RR), forget retention (FR), test retention (TR), where RR is given by,
$\text{RR}(f_{U},f_{R})=\frac{\text{RA}(\mathcal{D}_{R},f_{U})}{\text{RA}(\mathcal{D}_{R},f_{R})}\in[0,+\infty),$ (2)
and formulas for FR and TR can be derived using FA and TA, respectively. An unlearned model with a score of $1$ indicates that its accuracy perfectly matches the accuracy of the reference retrain model. A score below $1$ indicates that the model under-performs and a score above $1$ indicates that the model over-performs. We further define the retention deviation (RetDev) as:
$\text{RetDev}=|RR(f_{U},f_{R})-1|+|FR(f_{U},f_{R})-1|+|TR(f_{U},f_{R})-1|\in[0,+\infty),$ (3)
which provides information on the cumulative divergence of the unlearned model in terms of retention score. The closer to 0, the better as 0 indicates that the model perfectly matches the performance of the retrained model.
(3) Efficiency. run-time efficiency (RTE) of a MU method should ideally be lower than the naive approach of just retraining the model from scratch on the retain set. As the high computational cost of the retraining algorithm motivated the development of approximate MU methods, we evaluate how much faster each MU method is compared to the retraining algorithm. We define RTE of $U$ as the number of seconds it takes to complete, denoted as $\text{RT}(U)$ (we use the same machine and resources for all the experiments). To indicate the relative speedup compared to the retrained model, we define RTE of an unlearn method $U$ as:
$\frac{\text{RT}(U_{R})}{\text{RT}(U)}\in[0,+\infty),$ (4)
where $U_{R}$ denotes the retrain method. The RTE of retraining from scratch would thus be 1; any method with an RTE less than 1 is slower than retraining from scratch, and vice versa.
3 Machine Unlearning Methods
We briefly discuss the main unlearning methods considered.
3.1 Classical Baselines
FineTune (FT) finetunes the original model $f_{O}$ on only the retain set $\mathcal{D}_{R}$ for several epochs.
Successive Random Labels (SRL) the model is trained on both the forget set $\mathcal{D}_{F}$ and $\mathcal{D}_{R}$ where the labels of $\mathcal{D}_{F}$ are randomly assigned at each epoch.
Gradient Ascent (GA) trains the model using gradient ascent steps on the $\mathcal{D}_{F}$.
3.2 State-of-the-art MU methods
Expanding upon the classical baselines, we additionally evaluate 15 recent MU methods. We first discuss the seven top-performing methods from the Machine Unlearning Competition 2023 on Kaggle.
Forget-Contrast-Strengthen (FCS) [1] minimizes the Kullback-Leibler Divergence (KLD) between the model’s output on $\mathcal{D}_{F}$ and a uniform distribution over the output classes, then alternatively optimizes a contrastive loss between the model’s outputs on $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$, and minimizes the cross-entropy loss on $\mathcal{D}_{R}$.
Masked-Small-Gradients (MSG)[2] accumulates gradients via gradient descent on the $\mathcal{D}_{R}$ and gradient ascent on the $\mathcal{D}_{F}$, then reinitialize weights with the smallest absolute gradients while dampening subsequent weights updates on the $\mathcal{D}_{R}$ for the other weights.
Confuse-Finetune-Weaken (CFW)[3] injects noise into the convolutional layers and then trains the model using a class-weighted cross-entropy on $\mathcal{D}_{R}$, then injects noise again toward the final epochs.
Prune-Reinitialize-Match-Quantize (PRMQ) [4] first prunes the model via L1 pruning, reinitializes parts of the model, optimises it using a combination of cross-entropy and a mean-squared-error on the entropy between the outputs of $f_{O}$ and $f_{U}$ on $\mathcal{D}_{R}$ and finally converts $f_{U}$’s weights to half-precision floats.
Convolution-Transpose [5] simply transposes the weights in the convolutional layers and trains on $\mathcal{D}_{R}$.
Knowledge-Distillation-Entropy (KDE) [6] uses a teacher-student setup. Both student and teacher start as copies of the original model, then the student’s first and last layers are re-initialised. The student $f_{U}$ minimizes its Kullback-Leibler Divergence (KLD) with the $f_{O}$ over $\mathcal{D}_{V}$, then minimizes a combination of losses: a soft cross-entropy loss between $f_{U}$ and $f_{O}$, a cross-entropy loss on outputs of $\mathcal{D}_{R}$ from $f_{U}$, and the KLD between $f_{U}$ and $f_{O}$ on $\mathcal{D}_{R}$.
Repeated-Noise-Injection (RNI) [7] first reinitialises the final layer of the model, then repeatedly injects noise in different layer of the model while training on the $\mathcal{D}_{R}$.
We further consider eight state-of-the-art methods introduced in the literature.
Fisher Forgetting (FF) [25, 20] adds noise to $f_{O}$ with zero mean and covariance determined by the 4th root of Fisher Information matrix with respect to $\theta_{O}$ on $\mathcal{D}_{R}$.
Influence Unlearning (IU)  [33, 47, 34] uses Influence Functions[18] to determine the change in $\theta_{O}$ if a training point is removed from the training loss. IU estimates the change in model parameters from $\theta_{O}$ to the model trained without a given data point. We use the first-order WoodFisher-based approximation from [34].
Catastrophic Forgetting - K (CF-K) [24] freezes the first layers then trains the last $k$ layers of the model on $\mathcal{D}_{R}$.
Exact Unlearning - K (EU-K) [24] freezes the first layers then restores the weights of the last $k$ layers to their initialization state. We randomly reinitialize the weights instead, so that the method no longer requires knowledge about the training process of $f_{O}$.
SCRUB [37] leverages a student-teacher setup where the model is optimised for three objectives: matching the teacher’s output distribution on $\mathcal{D}_{R}$, correctly predicting the $\mathcal{D}_{R}$ set and ensuring the output distributions of the teacher and student diverge on the $\mathcal{D}_{F}$
Saliency Unlearning (SaLUN) [20] determines via gradient ascent which weights of $\theta_{O}$ are the most relevant to $\mathcal{D}_{F}$, then trains the model simultaneously on $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$ with random labels on $\mathcal{D}_{F}$, while dampening the gradient propagation based on the selected weights.
Negative Gradient Plus (NG+)  [37] is an extension of the Gradient Ascent approach where additionally a gradient descent step is taken over the $\mathcal{D}_{R}$.
Bad Teacher (BT) [16] uses a teacher-student approach with two teachers: the original model, and a randomly initialized model - the bad teacher-, the student starts as copy of $f_{U}$ then learns to mimic the $f_{O}$ on $\mathcal{D}_{R}$ and the bad teacher on the $\mathcal{D}_{F}$.
4 Experimental Evaluation
[…]
The performance of the MU methods can change across datasets, model configurations, and model initializations; a reliable MU method remains consistent across these changes. For each method, dataset and model combination, we unlearn from Original models initialized using 10 different seeds and consider the average performance across seeds.
A further observation is that prior research tends to compare MU methods with default hyperparameters, potentially leading to a less competitive performance of the method. To ensure that we get the best performance out of each method, we perform three hyperparameter sweeps to find the best set of hyperparameters for each method. To ensure a fair comparison, we use same number of searches for each method. Each hyper-parameter sweep uses 100 trials to minimize four loss functions: Retain Loss ($\mathcal{L}_{\text{Retain}}$), Forget Loss ($\mathcal{L}_{\text{Forget}}$), Val Loss ($\mathcal{L}_{Val}$), and Val MIA ($\mathcal{L}_{\text{Val-MIA}}$) given by
$\mathcal{L}_{Retain}=\alpha\times|\text{RA}(f_{U})-\text{RA}(f_{R})|,\;\mathcal{L}_{Forget}=\beta\times|\text{FA}(f_{U})-\text{FA}(f_{R})|,$ (5)
$\mathcal{L}_{Val}=\gamma\times|\text{VA}(f_{U})-\text{VA}(f_{R})|,\;\mathcal{L}_{\text{Val-MIA}}=\eta\times\text{Disc}(\mathcal{D}_{V},f_{U}),$ (6)
where the $\mathcal{L}_{\text{Retain}}$ captures the divergence in accuracy between the retrained and unlearned model over the $\mathcal{D}_{R}$, $\mathcal{L}_{\text{Forget}}$ and $\mathcal{L}_{Val}$ capture the divergence over $\mathcal{D}_{F},\mathcal{D}_{V}$ respectively and $\mathcal{L}_{\text{Val-MIA}}$ captures whether the loss distributions over $\mathcal{D}_{F}$ and $\mathcal{D}_{V}$ are distinguishable from one another via the discernibility score defined in Section 2. We set $\alpha=\beta=\gamma=\frac{1}{3}$ and $\eta=1$ as we found these values to balance the importance of importance retention and the resilience to Membership Inference Attacks. Per unlearn method, we use the hyperparameter configuration that minimises the four loss terms when evaluating the method. Thus, for each unlearning method, we first unlearn $300$ models to do the hyper-parameter sweep, then unlearn $10$ models with the best set of hyper-parameters, leading to $5,580$ per dataset for a given architecture, leading to a total of $50,220$ for the 9 dataset / model combinations.
Ranking. A challenge in the comparison of MU method performance comes from the potential proximity of the evaluation metrics. As a simple example, suppose we have four methods $U_{1},...,U_{4}$ with accuracies: 98%, 99%, 50%, 1%, respectively; if we simply rank the methods, the rank itself would not be representative of the fact that e.g. $U_{1}$ and $U_{2}$ are much above $U_{3}$ and $U_{4}$. In order to enable distinctions based on proximities, we use Agglomerative Clustering and define cut-off points such that we obtain three clusters: (1) Best performers (G1), (2) Average performers (G2), and (3) Worst performers (G3). If a method does not produce 10 usable models, one per original model, it is assigned to a Failed group (F). For each method, we count the number of times it appears in each of the three groups (with nine being the maximum). To obtain a final ranking of the methods, we first rank the methods using the number of times it appears in the Best Performers group (G1); if ties occur, we use the Average Performers (G2) group to break them. If ties persist, the Worst Performers (G3) group serves as the final tie-breaker. This method ensures a clear and fair ranking by considering each performance group in order of importance.
5 Main Results
Table 1 presents the main results of our evaluations on MU methods based on Retention Deviation and Indiscernibility. The results for the run-time efficiency are shown in Table 2.
Table 1: Ranking by performance on Retention Deviation and Indiscernibility across datasets and architectures. We count the number of times each method appears in the Best Performers group (G1), Average performance group (G2) and Worst performers group (G3) (see §4). The final rank is computed based on the number of times the method appears in G1—with occurrences in G2 and G3 used to break ties if needed. If a method does not produce any usable models, it is assigned to a Failed group (F). Three methods appear in the top 3 for both performance measures: MSG (1st and 1st), CT (3rd and 1st) and KDE (3rd and 2nd).
Retention Deviation Indiscernibility
Rank Method G1 G2 G3 F Rank Method G1 G2 G3 F
1 FT 8 1 0 0 1 CT 9 0 0 0
1 MSG 8 1 0 0 1 MSG 9 0 0 0
2 PRMQ 7 2 0 0 2 CFW 7 2 0 0
3 CT 7 1 1 0 2 RNI 7 2 0 0
3 KDE 7 1 1 0 2 KDE 7 2 0 0
3 CFW 7 1 1 0 3 FT 6 3 0 0
4 FCS 6 3 0 0 3 PRMQ 6 3 0 0
4 SalUN 6 3 0 0 3 SalUN 6 3 0 0
5 NG+ 5 4 0 0 4 SRL 6 2 1 0
5 SRL 5 4 0 0 5 NG+ 5 4 0 0
6 SCRUB 4 3 1 1 5 FCS 5 4 0 0
7 BT 2 7 0 0 6 SCRUB 5 3 0 1
7 RNI 2 7 0 0 7 BT 4 5 0 0
8 CF-k 2 3 2 2 8 CF-k 1 2 4 2
9 IU 1 0 2 6 9 EU-k 1 2 2 4
10 EU-k 0 5 0 4 10 GA 0 4 4 1
11 GA 0 1 7 1 11 IU 0 0 3 6
12 FF 0 0 0 9 12 FF 0 0 0 9
On the reliability of baselines.
The commonly used baseline FT trains the original model only on the retain set for several epochs to enable the model to forget information about the forget set. In our evaluation, FT performs best based on the Retention Deviation, and is ranked third based on Indiscernibility. The latter observation may come from the fact that FT does not explicitly unlearn the forget set or perturb the model parameters. Based on these results we conclude it is a reasonable baseline to evaluate against. We however remark that since the mechanism underlying FT (training on the Retain set to maintain performance) is common to many other methods, these methods may inherit its susceptibility to MIA. Another common baseline, GA which performs gradient ascent on the forget set, performs poorly across both metrics. Its more recent variation NG+, which uses an additional retain set correction, ranks fifth for both metrics, making it a more suitable baseline.
Table 2: Run Time Efficiency on ResNet for the top performing methods. CT is the fastest on average, MSG runs up to 5x faster than naive retraining.
CIFAR-10 CIFAR-100 MNIST FashionMNIST UTKFace Average
Unlearner
MSG 6.80 4.49 4.29 3.32 7.57 5.29
CFW 4.67 6.17 4.29 4.90 5.54 5.11
PRMQ 4.93 4.34 3.77 3.77 5.88 4.54
CT 17.49 11.82 5.83 4.47 13.34 10.59
KDE 6.33 3.98 3.27 3.22 8.19 5.00
FT 8.15 5.16 5.07 4.29 5.78 5.69
On the reliability of newly proposed unlearning methods.
MSG obtains a first rank in both Performance Retention Deviation and Indiscernibility. Its unique approach identifies the parameters in the Convolutional layers that most contribute to the information to be forgotten. This strategy, differing from FT, allows MSG to retain performance from the Retain set while modifying the weights that are more relevant to the Forget set. PRMQ ranks second in performance retention, making it one of the top performers. However, it suffers from the same lower performance in Indiscernibility as FT. PRMQ however does not leverage the Forget set; instead, it performs a form of knowledge distillation by attempting to reproduce the results of the original models on the Retain set. Additionally, during the pruning phase, it reinitializes the weights for the MLP and Convolutional layers. CT ranks third in Performance Retention and first in Indiscernibility. It is interesting to note that CT and MSG are both consistently among the top performers in Indiscernibility, where FT performs poorly.
On the robustness across architectures.
Critical for a good MU method is its ability to generalise across various DNN architectures. We conducted experiments with both ResNet18 and TinyViT. Methods such as CT, despite being tailored to Convolutional Neural Networks (CNN) models, still perform competitively when applied to Vision Transformers. Methods such as CT and MSG, while proposed for CNN layers, work well on Vision Transformer as one can leverage the 2D Convolutions used in Positional Encoders. We provide additional details on the ranking based on architectures in Appendix F.
On the speed of the unlearn methods.
MSG, the best MU candidate runs  7.6x faster than retraining from scratch on UTKFace and  3.3x on FashionMNIST. On average MSG runs  5.3x faster than retraining. The fastest method is CT which achieves a speedup of  17.5x compared to Retraining. This speedup stems from its simple approach: transpose the weights of the convolutional layers.
On the performance when evaluated against a stronger MIA.
From the results above, we note that there exist MU methods that perform fast and reliably across datasets, architectures, and initializations. Specifically, CT, FT, MSG seem strong candidates for reliable MU methods. However, as has also been highlighted in prior work [46], MIA has been debated as a strong metric, as its ability to assess MU is hampered by its own ability to infer data membership. In line with this, recent works have introduced more powerful variations on MIA [51] with [37] proposing the stronger U-LiRA attack for MU. For the best performers in Table 1, we apply the attack setup from [31] where we generate a total of 640 models (with varying train, retain and forget sets) and for each unlearn method perform a hyperparameter sweep to find the best configuration (for details see Appendix D). We determine for each data point its U-LiRA Inference Accuracy and report for each method its average and standard deviation (see Figure 1). From this, we conclude that MSG as well as CT resist U-LiRA attacks. Both MSG and CT thus not only rank first in terms of Performance Retention and Indiscernibility based on U-MIA, but also are robust against a stronger variation of MIA.
Figure 1: U-LiRA on CIFAR-10 on ResNet models. Both CT and MSG, which ranked first against U-MIA, showed great resilience against the U-LiRA attack.
6 Discussion and Conclusion
The increasing focus on data privacy and trustworthiness of machine learning models underscores the need for robust and practical methods to unlearn and remove the influence of specific data from trained models. Due to the growing size of models, we require methods that avoid the computationally costly retraining from scratch. In this work we performed a comprehensive comparison of approximate unlearning methods across various models and datasets aimed to address this critical issue.
We experimentally compared 18 methods across different datasets and architectures, focusing on assessing the method’s ability to maintain privacy and accuracy while being computationally efficient and reliable across datasets, architectures and random seeds. Our findings indicate that Masked-Small-Gradients, which accumulates gradients via gradient descent on the data to remember and gradient ascent on the data to forget to determine which weights to update, consistently outperforms for all metrics across the studied datasets, architectures, and initialization seeds. Similarly, Convolution Transpose, which leverages the simple transposition in convolutional layers, performed strongly.
Both CT and MSG were resistant against both a population-based Membership Inference Attack (MIA) and a stronger, per-sample attack (U-LiRA). However, a core challenge of approximate unlearning is that these methods will only be as strong as the attacks against which they are tested. As stronger and more complex attacks emerge, some approximate unlearning methods might no longer be as efficient as initially expected. This highlights the need for continuous evaluation and adaptation of unlearning methods to maintain their effectiveness. We also conducted experiments based on L2 distances, but found that no method consistently got close to the reference models’ weights, we provide further information in Appendix G.
Limitations. Due to computational costs, we limited our analysis to Tiny Vision Transformers and ResNet; a further investigation of other architectures could provide useful insights. We did not investigate different amounts of unlearning samples, which some methods are known to be sensitive to [37]. We did not consider repeated deletion, instead we assume that there is a single forget set and that the unlearning process happens once, as is common in the literature, nonetheless, in practical applications one might need to unlearn different smaller forget sets over time and some unlearning methods might not work as well under such scenario. We finally remark once again on the difficulty of evaluation for approximate unlearning [31]: while these methods provide significant gains in efficiency, novel attacks might highlight yet unknown weaknesses of the unlearning processes.
Future work. First, we put our focus on natural image data, however, machine unlearning is relevant to other data types such as medical images or other modalities such as time series, audio and speech, or language data. Second, we focus on the classification task, however, other learning tasks would greatly benefit from machine unlearning too. For instance removing concepts from generative models for images [20] or poisoned data in language models [31]. Third, this work focuses on empirically benchmarking approximate machine unlearning methods. We do not provide a theoretical analysis of these methods or a rigorous comparison with exact unlearning algorithms.
Impact statement. This paper aims to highlight the importance of effectively assessing approximate machine unlearning methods. Our goal is to stress the need for evaluating new unlearning methods against more reliable baselines and experimental setups. Additionally, it is crucial to assess the consistency of a new unlearning method across various datasets and model architectures. Without such a thorough evaluations, proposed unlearning methods may provide a false sense of privacy and safety, ultimately limiting their effectiveness for data regulation.
7 Acknowledgements
Xavier F. Cadet is supported by UK Research and Innovation (UKRI Centre for Doctoral Training in AI for Healthcare grant number EP/S023283/1). Hamed Haddadi is supported by the EPSRC Open Plus Fellowship (EP/W005271/1: Securing the Next Billion Consumer Devices on the Edge).
8 Impact Statement
This paper presents work whose goal is to advance the fields of Machine Learning and Machine Unlearning. There are many potential societal consequences of our work. We aim to raise awareness on the necessity of developing robust methods and evaluating the consistency of Machine Unlearning methods across datasets, as relying on inconsistent Machine Unlearning mechanism could go against Data Regulations.
Appendix
The appendix is structured as follows:
  • Related Works (Section A) provides an overview of the previous works related to benchmarking MU.
  • Datasets (Section B) contains information about the datasets, data augmentations, and hyper-parameters used to train the Original and Retrained models.
  • Neural Network Architectures (Section C) provides information about ResNet18 and TinyViT used throughout our benchmark.
  • Privacy Evaluation (Section D) gives the descriptions of the U-MIA and U-LiRA evaluation metrics.
  • Per Dataset Results (Section E) shows some experimental results in terms of accuracy, retention, privacy metrics, and runtime efficiency per dataset for the 9 combinations of datasets and DNN architectures considered in our work.
  • Per Architectures Ranking (Section F) provides Performance Retention Deviation and Indiscernibility Rankings separated for ResNet18 and TinyViT.
  • L2 Distances between Model Weights (Section G) shows L2 distances computed between the Unlearned, Original, and Retrained models.
  • Requirements (Section H) which describes the compute resources.
Appendix A Related Works
A.1 Machine Unlearning
Machine Unlearning is often first associated with the work from Cao et al. [11], followed by Bourtoule et al. [9], which proposes SISA (Sharded, Isolated, Sliced, Aggregated) as an exact unlearning method. For a recent overview of MU, we refer to the survey from Xu et al. [50], which provides a taxonomy of common unlearning methods. Furthermore, Zhang et al. [52] review MU through privacy-preserving and security lenses. The authors cover the Confidentiality, Integrity, and Availability security triad and the need for Data Lineage, which relates to following the movement of data in a machine learning pipeline and understand from where it originates, where it is stored, and how it percolates in the system through transformation. Some might have information on common MU verification methods, privacy evaluation metrics, and datasets, we defer to the work of Nguyen et al. [39], and Shaik et al. [43].
In the following, we focus on the MU taxonomy from Xu et al. [50], which considers Data Reorganization and Model Manipulation:
(1) Data Reorganization methods focus on directly modifying the data to perform unlearning. It is divided into Data Obfuscation, Data Pruning, and Data Replacement.
Data Obfuscation refers to modifying the dataset to obscure the influence of the data to be unlearned: random relabeling and retraining [28], SRL (Successive Random Labels), and Saliency Unlearning (SalUN).
Data Pruning usually relies on dividing the dataset into multiple sub-datasets and training sub-models on these subsets. This is the category to which SISA [9] relates. Our work does not consider methods associated with this setting as they assume the training process.
Data Replacement attempts to unlearn by replacing the original dataset with transformed data that simplifies unlearning specific samples. For instance, Cao et al. [11], replace the training data with summations of efficiently computable transformations. Like data pruning, these methods tend to make strong assumptions about the training process.
(2) Model Manipulation methods directly adjust the model parameters to remove the influence of specific data points. Model manipulation is divided into Model Shifting, Model Replacement, and Model Pruning.
Model Shifting directly updates the model parameters to offset the influence of the unlearned samples, such as using a single step of Newton’s method on model parameters [30] or decremental updates [40], in our benchmark Fisher Forgetting (FF), Influence Unlearning (IU), and SalUN would represent these approaches.
Model Replacement uses pre-calculated parameters that do not reflect the data to forget to replace parts of the trained model. For instance, when using decision trees, one can replace nodes affected by the forget set by pre-calculated node [41].
These methods often make strong assumptions about the training process and the overall model.
Model Pruning prunes specific parameters from the trained models to remove the influence of certain samples [34] or Prune-Reinitialize-Match-Quantize (PRMQ) [4] which prunes the model via L1 pruning, reinitializes parts of the model then train the model on $\mathcal{D}_{R}$.
A.2 Machine Unlearning for Deep Neural Networks
Initially, MU research primarily focused on linear models such as linear regression and logistic models. Such models allow for the design of methods that assume the convexity of the loss function, rendering them less practical for DNN-based approaches. Since DNNs can memorize parts of their training data, they are particularly relevant targets for MU, even more so when they have been trained on large amounts of potentially personal data. For Deep Learning models, unlearning raises additional challenges: 1) the non-convexity of the loss function of Deep Neural Networks [15], 2) the size of the models inducing large computational costs, 3) the randomness coming from the model’s training process, such as the initialization seed, randomness in the mini-batch generation process, and 4) the fact that any model update impacts subsequent versions of the models, namely the weights at epoch $n+1$ directly depend on the weights at update $n$.
When considering MU for DNN, Xu et al. [50] notes that a standard scheme DNN is to focus only on the final layer, as it is expected for this layer to be the most relevant for the downstream task, and stems from the early works MU. Nonetheless, Goel et al. [24] showed that simply modifying the final layer is often insufficient to remove information related to $\mathcal{D}_{f}$. However, other approaches, such as those from Golatkar et al.  [25, 26, 27], attempt to unlearn the full model via methods derived from Information Theory. For instance, weight scrubbing on trained models can be done by approximating the Fisher information matrix.
A.3 Post-Hoc Machine Unlearning
While proactively designing deep-learning pipelines with built-in unlearning methods such as SISA can greatly simplify the unlearning process, many contemporary services relying on DNNs were not deployed with unlearning in mind. This motivates searching for methods that can unlearn from already trained models without making assumptions about the training process.
Thus, we focus on post-hoc MU, a scenario where we assume that the unlearning method is agnostic to the original training process of the model. Under such a scenario, differences exist in terms of data availability at unlearning time. For instance, whether one has access to the original training data $\mathcal{D}$, the retain set $\mathcal{D}_{R}$, the forget set $\mathcal{D}_{F}$, or even some external set such as the validation set $\mathcal{D}_{V}$. Therefore, careful consideration should be given to the data requirement associated with an unlearning method. Indeed, some might require having access to both $\mathcal{D}_{R}$ and $\mathcal{D}_{F}$ at the unlearning time, while others assume that $\mathcal{D}_{R}$ is no longer available [17] making them more practical in real-world scenarios. Throughout our benchmark, we make the same assumption as the NeurIPS2023 Unlearning Challenge [45], where the unlearning methods had access to $f_{O},\mathcal{D}_{R},\mathcal{D}_{F},\mathcal{D}_{V}$
A.4 Machine Unlearning and Differential Privacy
We based our Unlearning definition on Sekhari et al. [42] and refer to their work on the distinction between Differential Privacy and the objective of Machine Unlearning. Differential privacy, in a high-level picture, is a method for publicly sharing aggregated information about a population by describing the patterns discovered among the groups within the dataset while withholding specific information about individual data points. A randomized algorithm $\mathcal{A}$ is $(\varepsilon,\delta)$-differentially private if for all datasets $D_{1}$ and $D_{2}$ that differ on a single data point, and all $S\subseteq\text{Range}(\mathcal{A})$,
$\Pr[\mathcal{A}(D_{1})\in S]\leq e^{\varepsilon}\cdot\Pr[\mathcal{A}(D_{2})\in S]+\delta.$ (7)
In this definition, $\varepsilon$ (epsilon) is a non-negative parameter that measures the privacy loss, with smaller values indicating stronger privacy. The parameter $\delta$ represents the probability of breaking differential privacy, ideally close to or equal to 0.
Despite enabling provable error guarantees for Unlearning methods, Differential Privacy requires strong model and algorithmic assumptions, making MU, derived from it, potentially less effective against practical adversaries [34].
Appendix B Datasets
CIFAR-10
CIFAR-10 is a widely used dataset in computer vision and machine learning. It comprises 60,000 32x32 color images in 10 different classes, with 6,000 images per class. The dataset is divided into 50,000 training images and 10,000 testing images. CIFAR-10 represents a diverse range of everyday objects, such as airplanes, automobiles, birds, and cats, making it a challenging task for image classification. The simplicity of the images combined with the variety of categories makes CIFAR-10 a suitable dataset to test the efficacy of machine unlearning algorithms in effectively unlearning information without compromising the model’s performance on the remaining data.
Data Augmentations: random cropping to 32x32 with 4-pixel padding, 50% random horizontal flipping, and per-channel normalization with a mean of $[0.4919,0.4822,0.4465]$ and standard deviation of $[0.2023,0.1994,0.2010]$. At test time, we resize to 32x32 and normalize.
CIFAR-100
CIFAR-100 is a more complex extension of CIFAR-10, containing 100 classes with 600 images per class, split into 500 training images and 100 testing images per class. Each class is labeled with a "fine" label and grouped into 20 "coarse" labels, adding another layer of classification difficulty. The increased number of classes and finer granularity make CIFAR-100 an intriguing dataset for machine unlearning benchmarks. It poses a more significant challenge for models to forget specific classes or groups while retaining knowledge of others, thus testing the unlearning algorithms’ precision and effectiveness in handling more granular and complex datasets.
Data Augmentations: random cropping to 32x32 with 4-pixel padding, 50% random horizontal flipping, and per-channel normalization with a mean of $[0.5071,0.4865,0.4409]$ and standard deviation of $[0.2673,0.2564,0.2762]$. At test time, we resize to 32x32 and normalize.
MNIST
The MNIST dataset is a well-known benchmark in handwritten digit recognition. It comprises 70,000 grayscale images of handwritten digits (0-9), 60,000 used for training, and 10,000 for testing. Each image is 28x28 pixels in size. We consider MNIST due to its simplicity and extensive research and development history. The simplicity of MNIST allows researchers to focus on the fundamental aspects of unlearning techniques without the additional complexity introduced by color or high resolution, providing a clear assessment of the effectiveness of unlearning algorithms in a controlled setting.
Data Augmentations: conversion to 3 channels, resizing to 32x32 such that both ResNet18 and TinyViT use the same input resolution, 50% random horizontal flipping, and per-channel normalization with a mean of $[0.1307,0.1307,0.1307]$ and standard deviation of $[0.3081,0.3081,0.3081]$. We convert to 3 channels at test time, resize to 32x32, and normalize.
Fashion MNIST
Fashion MNIST is a more challenging replacement for MNIST. It contains 70,000 grayscale images of fashion items in 10 categories: shirts, trousers, and sneakers. Like MNIST, each image is 28x28 pixels, but the increased complexity and variability of clothing items make it a more challenging classification task. Fashion MNIST provides a more realistic and intricate dataset than MNIST, testing the unlearning algorithms’ ability to handle real-world-like variability and ensuring that they can effectively remove learned information while maintaining performance on a moderately complex dataset.
Data Augmentations: conversion to 3 channels, resizing to 32x32, 50% random horizontal flipping, and per-channel normalization with a mean of $[0.2860,0.2860,0.2860]$ and standard deviation of $[0.3560,0.3560,0.3560]$. We convert to 3 channels at test time, resize to 32x32, and normalize.
UTKFace
UTKFace is a large-scale face dataset containing over 20,000 images of faces with annotations of age, gender, and ethnicity. The images vary in size and cover a wide range of ages, from 0 to 116. UTKFace is particularly interesting due to the sensitive nature of the data and the need for privacy-preserving techniques.
Data Augmentations: resizing to 224x224, and per-channel normalization with a mean of $[0.485,0.456,0.406]$ and standard deviation of $[0.229,0.224,0.225]$. We apply the same transformation at test time.
For each dataset, the Original and Retrained models are trained using the same hyper-parameters (provided in Table 3)
Table 3: Summary of the number epochs, learning rate, and batch size for each dataset and model used to train the Original and Retrained models.
Dataset Model Epochs Learning Rate Batch Size
FashionMNIST ResNet18 50 0.1 256
TinyViT 50 0.1 256
MNIST ResNet18 50 0.1 256
TinyViT 50 0.1 256
CIFAR-10 ResNet18 182 0.1 256
(LiRA) ResNet18 91 0.1 256
TinyViT 182 0.1 256
CIFAR-100 ResNet18 182 0.1 256
TinyViT 182 0.1 256
UTKFace ResNet18 50 0.1 128
TinyViT 50 0.1 128
Appendix C Neural Network Architectures
We consider two families, ResNet (Residual Network) [32] and ViT (Vision Transformer) [19], which are prominent architectures in computer vision. We consider ResNet18 and a TinyViT[48] with approximately 11M learnable parameters for a fair comparison between two fundamentally different architectures. This provides insights into how architectural differences impact the unlearning process and helps understand the trade-offs between convolutional and transformer-based models regarding reliability and computational efficiency.
ResNet: ResNet18
Introduced by He et al. [32], it facilitates the training of deep networks through shortcut connections, which mitigates the problem of vanishing gradients. The ResNet18 is known for its balance between performance and computational efficiency.
ViT: TinyViT
Vision Transformer (ViT), introduced by Dosovitskiy et al. [19], adapts the transformer architecture to image classification by treating images as sequences of patches. We consider TinyViT from Wu et al. [48], as it is a compact version of ViT designed to be parameter-efficient while maintaining high performance.
Appendix D Privacy Evaluation
D.1 Unlearning-Membership Inference Attack (U-MIA)
A common approach to evaluate the quality of unlearning methods is to attack the unlearned models with a form of Membership inference Attack (MIA). Membership Inference Attacks attempt to determine whether a specific data point was part of the model train data. The efficacy of the Membership Inference Attack has been used as a metric to evaluate the success of unlearning algorithms. A general approach to such an attack is as follows. Assume $f\theta$ is a trained model with parameters $\theta$, and let $\mathcal{L}$ be a loss function, such as the cross-entropy loss. Then, compute the losses for each sample from two sets of data $A$ and $B$ (of equal size) and train a binary classification model such as logistic regression with labels $y^{A}_{i}=1$ for points $i$ in $A$ and $y^{B}_{i}=0$ for points $i$ in $B$. An accuracy score from the classifier close to $1.0$ indicates that the classifier can perfectly distinguish between samples from $A$ and $B$ based on the loss values. A score of $0.5$ indicates that the ability to distinguish is close to random.
D.2 Unlearning -Likelihood Ratio Attack (U-LiRA)
The performance of a general MIA can be improved by considering, e.g., a per-sample attack such as LiRA [13, 31]. For any given point, we wish to determine whether the outputs from the unlearned models differ from those of models that have never seen the data point. To assess the attack robustly, we evaluate it across multiple models, using shadow models trained on various retain/forget sets. Specifically, we first train $n$ models based on $n$ splits of the training data. This train data is then split into 10 random retain and forget splits, and hence, we unlearn a total of $10n$ models. We then perform hyper-parameter sweeps, similar to what we do in the original results and unlearn using the optimal hyper-parameters, except that we consider $\frac{n}{2}$ sweeps and conduct $200$ trials per sweep to determine the best hyper-parameters. In our setting, we set $n=64$.
Appendix E Per dataset results
Here, we present the results for both ResNet18 and the TinyViT across datasets.
ResNet18 We provide the tables with Retain Accuracy (RA), Forget Accuracy (FA), Test Accuracy (TA), Retain Retention (RR), Forget Retention (FR), Test Retention (TR), Performance Retention Deviation (RetDev), Indiscinerbility concerning the Test Set (Indisc), U-MIA on the Test set (T-MIA) and RunTime Efficiency (RTE) for every dataset using the ResNet18 model on MNIST (Table 4), FashionMNIST (Table 5), CIFAR-10 (Table 6), CIFAR-100 (Table 7) and UTKFace (Table 8). In general, CIFAR-100 provides the most visible differences, as the performance on the retain set is much higher than on the test. Datasets such as MNIST and FashionMNIST tend to show smaller differences between the methods as the performance on both the Retain and Test sets are similar, to begin with.
Table 4: MNIST - ResNet18
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
unlearner
BT 1.00 1.00 0.99 1.00 1.01 1.00 0.01 0.99 0.50 9.76
CF-k 1.00 1.00 0.99 1.00 1.01 1.00 0.01 0.98 0.51 5.86
CFW 1.00 1.00 0.99 1.00 1.00 1.00 0.00 1.00 0.50 4.29
CT 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 5.83
EU-k - - - - - - - - - -
FCS 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 2.11
FF - - - - - - - - - -
FT 1.00 0.99 0.99 1.00 1.00 1.00 0.00 0.99 0.51 5.07
GA 0.98 0.98 0.97 0.98 0.99 0.98 0.04 0.99 0.51 33.97
IU - - - - - - - - - -
KDE 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 3.27
MSG 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 4.29
NG+ 1.00 0.99 0.99 1.00 1.00 1.00 0.01 1.00 0.50 3.12
O 1.00 1.00 0.99 1.00 1.01 1.00 0.01 0.98 0.51 1.10
PRMQ 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 3.77
R 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 1.00
RNI 1.00 1.00 1.00 1.00 1.00 1.00 0.01 1.00 0.50 4.70
SCRUB - - - - - - - - - -
SRL 1.00 1.00 0.99 1.00 1.01 0.99 0.01 0.99 0.51 8.55
SalUN 1.00 1.00 0.99 1.00 1.00 0.99 0.01 0.99 0.50 3.19
Table 5: FashionMNIST - ResNet18
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 1.00 0.96 0.92 1.00 1.03 0.99 0.04 0.98 0.51 13.57
CF-k 0.98 0.97 0.91 0.98 1.05 0.98 0.09 0.92 0.54 16.31
CFW 1.00 0.95 0.92 1.00 1.02 0.99 0.03 0.97 0.51 4.90
CT 1.00 0.92 0.92 1.00 0.99 0.99 0.02 0.99 0.50 4.47
EU-k - - - - - - - - - -
FCS 0.98 0.93 0.91 0.98 1.00 0.98 0.04 0.98 0.51 2.79
FF - - - - - - - - - -
FT 1.00 0.95 0.92 1.00 1.02 1.00 0.02 0.98 0.51 4.29
GA - - - - - - - - - -
IU 1.00 1.00 0.93 1.00 1.08 1.00 0.08 0.87 0.56 14.97
KDE 1.00 0.93 0.92 1.00 1.00 1.00 0.01 0.99 0.50 3.22
MSG 1.00 0.93 0.91 1.00 1.00 0.99 0.01 0.98 0.51 3.32
NG+ 0.99 0.94 0.91 0.99 1.01 0.99 0.03 0.99 0.49 3.21
O 1.00 1.00 0.93 1.00 1.08 1.00 0.08 0.87 0.56 1.11
PRMQ 0.98 0.93 0.91 0.98 1.00 0.99 0.04 0.98 0.51 3.77
R 1.00 0.93 0.92 1.00 1.00 1.00 0.00 1.00 0.50 1.00
RNI 0.98 0.93 0.91 0.98 1.00 0.98 0.04 0.98 0.51 3.25
SCRUB 0.95 0.93 0.90 0.95 1.00 0.98 0.08 0.97 0.51 5.62
SRL 1.00 0.97 0.92 1.00 1.04 1.00 0.05 0.98 0.51 24.85
SalUN 0.99 0.97 0.92 0.99 1.04 0.99 0.06 0.98 0.51 25.02
Table 6: CIFAR-10 - ResNet18
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 0.94 0.87 0.84 0.94 1.00 0.96 0.10 0.97 0.48 51.96
CF-k - - - - - - - - - -
CFW 1.00 0.81 0.80 1.00 0.92 0.92 0.16 1.00 0.50 4.67
CT 1.00 0.82 0.81 1.00 0.93 0.93 0.14 0.99 0.50 17.49
EU-k - - - - - - - - - -
FCS 0.99 0.86 0.84 0.99 0.98 0.96 0.07 0.98 0.49 22.53
FF - - - - - - - - - -
FT 1.00 0.84 0.82 1.00 0.96 0.95 0.09 1.00 0.50 8.15
GA 0.91 0.89 0.81 0.91 1.02 0.93 0.18 0.92 0.54 91.48
IU 0.95 0.94 0.84 0.95 1.08 0.97 0.16 0.91 0.55 64.61
KDE 0.98 0.84 0.80 0.98 0.96 0.92 0.15 0.97 0.52 6.33
MSG 1.00 0.85 0.83 1.00 0.97 0.95 0.08 0.99 0.51 6.80
NG+ 0.97 0.89 0.85 0.98 1.02 0.97 0.07 0.98 0.51 12.89
O 0.96 0.96 0.85 0.96 1.10 0.98 0.16 0.89 0.55 1.08
PRMQ 1.00 0.86 0.83 1.00 0.98 0.95 0.07 0.98 0.51 4.93
R 1.00 0.87 0.87 1.00 1.00 1.00 0.00 1.00 0.50 1.00
RNI 1.00 0.83 0.81 1.00 0.95 0.93 0.12 0.99 0.50 3.60
SCRUB 0.99 0.85 0.85 0.99 0.97 0.98 0.07 0.99 0.50 2.57
SRL 0.99 0.93 0.84 0.99 1.06 0.97 0.10 0.98 0.49 5.52
SalUN 0.98 0.90 0.84 0.98 1.04 0.97 0.08 0.96 0.48 18.04
Table 7: CIFAR-100 - ResNet18. CIFAR-100 provides the most visible comparison as there is a large gap in performance between the Retain Set and Test set, this leads to much larger RetDev scores.
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 0.98 0.68 0.54 0.98 1.25 0.99 0.27 0.95 0.48 9.39
CF-k 1.00 0.83 0.56 1.00 1.53 1.02 0.55 0.73 0.63 5.91
CFW 0.98 0.43 0.43 0.98 0.79 0.78 0.44 1.00 0.50 6.17
CT 0.99 0.53 0.53 0.99 0.97 0.97 0.07 0.99 0.49 11.82
EU-k - - - - - - - - - -
FCS 0.98 0.54 0.55 0.98 0.99 1.01 0.04 0.92 0.54 3.02
FF - - - - - - - - - -
FT 0.98 0.55 0.54 0.98 1.02 0.98 0.05 0.99 0.50 5.16
GA 0.34 0.33 0.24 0.34 0.60 0.44 1.61 0.90 0.55 39.97
IU - - - - - - - - - -
KDE 0.99 0.52 0.51 0.99 0.95 0.94 0.11 0.99 0.50 3.98
MSG 0.91 0.38 0.38 0.91 0.69 0.69 0.71 1.00 0.50 4.49
NG+ 0.89 0.59 0.49 0.89 1.08 0.89 0.29 0.98 0.49 12.14
O 0.98 0.98 0.56 0.98 1.81 1.02 0.85 0.53 0.73 1.10
PRMQ 0.97 0.47 0.46 0.97 0.86 0.85 0.32 1.00 0.50 4.34
R 1.00 0.55 0.55 1.00 1.00 1.00 0.00 0.99 0.49 1.00
RNI 0.99 0.45 0.45 0.99 0.83 0.82 0.36 0.98 0.49 3.65
SCRUB 0.97 0.50 0.53 0.97 0.91 0.96 0.15 0.98 0.51 3.81
SRL 1.00 0.55 0.52 1.00 1.00 0.95 0.06 0.98 0.49 3.67
SalUN 0.98 0.49 0.51 0.98 0.91 0.93 0.18 0.99 0.49 10.66
Table 8: UTKFace - ResNet18
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 1.00 0.74 0.73 1.00 1.00 0.96 0.04 0.99 0.50 12.48
CF-k 1.00 1.00 0.75 1.00 1.34 0.99 0.35 0.70 0.65 5.35
CFW 1.00 0.76 0.76 1.00 1.02 1.00 0.02 1.00 0.50 5.54
CT 1.00 0.75 0.76 1.00 1.01 1.00 0.01 0.99 0.50 13.34
EU-k 0.72 0.61 0.59 0.72 0.82 0.77 0.68 0.99 0.51 11.42
FCS 0.90 0.70 0.70 0.91 0.94 0.93 0.23 0.99 0.50 4.33
FF - - - - - - - - - -
FT 1.00 0.76 0.77 1.00 1.02 1.01 0.04 1.00 0.50 5.78
GA 0.49 0.47 0.40 0.49 0.63 0.53 1.34 0.92 0.54 235.10
IU 1.00 1.00 0.76 1.00 1.34 1.01 0.35 0.62 0.69 33.77
KDE 0.99 0.79 0.76 0.99 1.06 1.00 0.07 0.97 0.52 8.19
MSG 1.00 0.80 0.76 1.00 1.08 1.00 0.08 0.96 0.52 7.57
NG+ 0.94 0.80 0.72 0.95 1.07 0.95 0.18 0.99 0.51 6.73
O 1.00 1.00 0.76 1.00 1.34 1.01 0.35 0.61 0.69 1.09
PRMQ 0.91 0.72 0.72 0.91 0.97 0.95 0.17 1.00 0.50 5.88
R 1.00 0.75 0.76 1.00 1.00 1.00 0.00 1.00 0.50 1.00
RNI 0.96 0.75 0.73 0.96 1.01 0.96 0.08 0.98 0.51 5.11
SCRUB 0.80 0.76 0.69 0.80 1.01 0.92 0.29 0.94 0.53 4.64
SRL 1.00 0.80 0.73 1.00 1.08 0.97 0.11 0.99 0.51 12.05
SalUN 0.97 0.79 0.73 0.98 1.06 0.96 0.12 0.97 0.52 36.80
TinyViT We provide the tables with RA, FA, TA, RR, FR, TR, RetDev, Indisc, T-MIA and RTE for MNIST (Table 9), FashionMNIST (Table 10), CIFAR-10 (Table 11) and CIFAR-100 (Table 12) using the TinyViT model.
Table 9: MNIST - TinyViT
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 1.00 1.00 0.99 1.00 1.01 1.00 0.01 1.00 0.50 4.39
CF-k 1.00 1.00 0.99 1.00 1.01 1.00 0.01 0.99 0.51 123.88
CFW 1.00 0.99 0.99 1.00 1.00 1.00 0.00 0.99 0.50 5.15
CT 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 5.92
EU-k 1.00 1.00 0.99 1.00 1.01 1.00 0.01 0.99 0.50 11.93
FCS 1.00 1.00 0.99 1.00 1.01 1.00 0.01 0.99 0.49 7.42
FF - - - - - - - - - -
FT 1.00 0.99 0.99 1.00 1.00 1.00 0.01 1.00 0.50 7.69
GA 0.97 0.97 0.96 0.97 0.98 0.97 0.08 0.99 0.51 390.99
IU - - - - - - - - - -
KDE 1.00 0.99 0.99 1.00 1.00 1.00 0.01 1.00 0.50 5.35
MSG 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 5.21
NG+ 1.00 0.99 0.99 1.00 1.00 1.00 0.00 0.99 0.50 2.74
O 1.00 1.00 0.99 1.00 1.01 1.00 0.01 0.99 0.51 0.97
PRMQ 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 6.88
R 1.00 0.99 0.99 1.00 1.00 1.00 0.00 1.00 0.50 1.00
RNI 1.00 0.99 0.99 1.00 1.00 1.00 0.01 1.00 0.50 8.03
SCRUB 0.99 0.99 0.99 0.99 1.00 1.00 0.01 1.00 0.50 3.54
SRL 1.00 0.99 0.99 1.00 1.00 1.00 0.01 0.98 0.51 16.37
SalUN 1.00 1.00 0.99 1.00 1.01 0.99 0.01 0.99 0.50 43.67
Table 10: FashionMNIST - TinyViT
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 0.97 0.94 0.91 0.97 1.01 0.99 0.04 0.98 0.51 3.48
CF-k - - - - - - - - - -
CFW 0.99 0.94 0.92 0.99 1.01 1.00 0.02 0.99 0.51 5.40
CT 0.98 0.92 0.91 0.98 0.99 0.99 0.04 0.99 0.50 6.00
EU-k 0.95 0.94 0.91 0.95 1.01 0.99 0.07 0.97 0.51 5.34
FCS 0.98 0.93 0.91 0.98 1.01 0.99 0.04 0.98 0.51 4.58
FF - - - - - - - - - -
FT 0.99 0.94 0.92 1.00 1.01 1.00 0.02 0.98 0.51 5.12
GA 0.92 0.91 0.85 0.92 0.99 0.93 0.17 0.93 0.53 50.72
IU - - - - - - - - - -
KDE 1.00 0.94 0.92 1.00 1.02 1.00 0.02 0.98 0.51 3.38
MSG 0.96 0.92 0.91 0.96 1.00 0.99 0.05 0.99 0.51 8.21
NG+ 0.97 0.92 0.91 0.97 1.00 0.99 0.05 0.98 0.51 13.65
O 1.00 1.00 0.92 1.00 1.08 1.00 0.08 0.89 0.56 0.97
PRMQ 0.98 0.94 0.91 0.98 1.01 0.99 0.04 0.98 0.51 4.67
R 1.00 0.93 0.92 1.00 1.00 1.00 0.00 1.00 0.50 1.00
RNI 0.97 0.94 0.91 0.97 1.01 0.99 0.05 0.97 0.51 5.00
SCRUB 0.96 0.95 0.91 0.96 1.03 0.99 0.08 0.96 0.52 9.56
SRL 0.98 0.94 0.91 0.98 1.01 0.99 0.04 1.00 0.50 9.41
SalUN 0.97 0.94 0.91 0.97 1.01 0.99 0.05 0.99 0.51 6.53
Table 11: CIFAR-10 - TinyViT
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 0.91 0.91 0.85 0.91 1.02 0.97 0.14 0.99 0.50 4.13
CF-k 0.99 0.89 0.84 0.99 1.00 0.95 0.06 0.96 0.52 5.18
CFW 0.98 0.87 0.84 0.99 0.98 0.96 0.07 0.98 0.51 28.85
CT 0.98 0.82 0.81 0.98 0.93 0.92 0.17 1.00 0.50 23.48
EU-k 0.90 0.90 0.84 0.90 1.02 0.95 0.16 0.97 0.52 43.25
FCS 0.98 0.84 0.83 0.98 0.95 0.94 0.13 0.99 0.49 5.15
FF - - - - - - - - - -
FT 1.00 0.87 0.84 1.00 0.98 0.95 0.07 0.98 0.51 5.85
GA 0.85 0.85 0.80 0.85 0.96 0.91 0.29 0.97 0.52 514.77
IU - - - - - - - - - -
KDE 0.97 0.86 0.84 0.97 0.97 0.96 0.11 0.99 0.50 5.27
MSG 1.00 0.85 0.83 1.00 0.96 0.94 0.10 0.99 0.51 7.38
NG+ 0.93 0.86 0.85 0.93 0.97 0.96 0.14 0.99 0.50 4.10
O 0.92 0.92 0.86 0.92 1.04 0.97 0.15 0.95 0.53 0.97
PRMQ 1.00 0.87 0.84 1.00 0.98 0.95 0.07 0.99 0.51 4.00
R 1.00 0.89 0.88 1.00 1.00 1.00 0.00 1.00 0.50 1.00
RNI 0.97 0.84 0.81 0.98 0.95 0.92 0.15 0.98 0.51 6.50
SCRUB 1.00 0.84 0.84 1.00 0.95 0.95 0.10 0.99 0.50 -
SRL 0.97 0.88 0.84 0.97 0.99 0.96 0.08 0.99 0.49 8.52
SalUN 0.96 0.89 0.85 0.96 1.00 0.96 0.08 0.99 0.50 8.30
Table 12: CIFAR-100 - TinyViT
RA FA TA RR FR TR RetDev Indisc T-MIA RTE
Unlearner
BT 0.82 0.66 0.57 0.82 1.11 0.96 0.33 0.94 0.53 31.63
CF-k 0.24 0.18 0.18 0.24 0.31 0.30 2.15 0.98 0.51 5.97
CFW 0.98 0.58 0.56 0.98 0.97 0.95 0.10 0.99 0.51 9.18
CT 0.97 0.55 0.55 0.98 0.93 0.93 0.17 0.99 0.49 9.80
EU-k 0.61 0.60 0.49 0.61 1.01 0.81 0.59 0.90 0.55 19.29
FCS 0.89 0.60 0.58 0.89 1.01 0.98 0.13 0.97 0.48 7.05
FF - - - - - - - - - -
FT 1.00 0.56 0.55 1.00 0.94 0.91 0.15 1.00 0.50 5.90
GA 0.60 0.58 0.46 0.60 0.97 0.77 0.66 0.87 0.56 91.91
IU - - - - - - - - - -
KDE 0.93 0.58 0.57 0.94 0.98 0.96 0.13 0.99 0.50 4.03
MSG 0.97 0.57 0.56 0.97 0.95 0.93 0.15 1.00 0.50 5.89
NG+ 0.84 0.57 0.55 0.84 0.95 0.92 0.29 0.96 0.52 3.01
O 0.87 0.87 0.61 0.87 1.46 1.02 0.61 0.74 0.63 0.98
PRMQ 0.95 0.62 0.57 0.95 1.03 0.96 0.13 0.95 0.52 5.58
R 1.00 0.60 0.60 1.00 1.00 1.00 0.00 1.00 0.50 1.00
RNI 0.85 0.52 0.51 0.85 0.87 0.85 0.42 0.99 0.50 5.38
SCRUB 0.77 0.64 0.57 0.77 1.08 0.96 0.35 0.93 0.53 6.17
SRL 0.98 0.57 0.57 0.98 0.96 0.96 0.11 0.97 0.49 5.92
SalUN 0.97 0.58 0.57 0.97 0.97 0.96 0.10 0.98 0.49 7.03
Appendix F Per architectures rankings
Here, we present the rankings across datasets for ResNet18 (Table 13) and TinyVit (Table 14). We note that some methods, such as RNI or NG+, are less efficient on the ViT architectures regarding Indiscernibility. However, methods such as SCRUB are less efficient regarding Retention Deviation on the ViT architecture.
Table 13: Ranking on ResNet
Retention Deviation Indiscernibility
Rank Method G1 G2 G3 F Rank Method G1 G2 G3 F
1 FT 5 0 0 0 1 CFW 5 0 0 0
2 FCS 4 1 0 0 1 CT 5 0 0 0
2 MSG 4 1 0 0 1 MSG 5 0 0 0
3 CT 4 0 1 0 1 RNI 5 0 0 0
3 KDE 4 0 1 0 2 FT 4 1 0 0
4 NG+ 3 2 0 0 2 KDE 4 1 0 0
4 PRMQ 3 2 0 0 2 NG+ 4 1 0 0
4 SalUN 3 2 0 0 2 PRMQ 4 1 0 0
5 CFW 3 1 1 0 3 FCS 3 2 0 0
6 SCRUB 3 0 1 1 3 SRL 3 2 0 0
7 SRL 2 3 0 0 3 SalUN 3 2 0 0
8 BT 1 4 0 0 4 SCRUB 3 1 0 1
8 RNI 1 4 0 0 5 BT 2 3 0 0
9 CF-k 1 2 1 1 6 EU-k 1 0 0 4
10 IU 1 0 2 2 7 GA 0 3 1 1
11 EU-k 0 1 0 4 8 CF-k 0 1 3 1
12 GA 0 0 4 1 9 IU 0 0 3 2
13 FF 0 0 0 5 10 FF 0 0 0 5
Table 14: Ranking of ViT
Retention Deviation Indiscernibility
Rank Method G1 G2 G3 F Rank Method G1 G2 G3 F
1 CFW 4 0 0 0 1 CT 4 0 0 0
1 MSG 4 0 0 0 1 MSG 4 0 0 0
1 PRMQ 4 0 0 0 2 KDE 3 1 0 0
2 CT 3 1 0 0 2 SalUN 3 1 0 0
2 FT 3 1 0 0 3 SRL 3 0 1 0
2 KDE 3 1 0 0 4 BT 2 2 0 0
2 SRL 3 1 0 0 4 CFW 2 2 0 0
2 SalUN 3 1 0 0 4 FCS 2 2 0 0
3 FCS 2 2 0 0 4 FT 2 2 0 0
3 NG+ 2 2 0 0 4 PRMQ 2 2 0 0
4 BT 1 3 0 0 4 RNI 2 2 0 0
4 RNI 1 3 0 0 4 SCRUB 2 2 0 0
4 SCRUB 1 3 0 0 5 NG+ 1 3 0 0
5 CF-k 1 1 1 1 6 CF-k 1 1 1 1
6 EU-k 0 4 0 0 7 EU-k 0 2 2 0
7 GA 0 1 3 0 8 GA 0 1 3 0
8 FF 0 0 0 4 9 FF 0 0 0 4
8 IU 0 0 0 4 9 IU 0 0 0 4
Appendix G L2 Distances between model weights
The distance between the Unlearned and Retrained models has also been considered in the literature to evaluate MU. Nevertheless, we observe that models end up at a similar distance to the Retrained model, with significant differences in performance. We further note that one challenging aspect of the L2 distance comparison is the different factors of Weight Decay used by the MU method. The hyper-parameter searches determine these Weight Decay factors, which can significantly vary from one unlearning method to another, making it challenging to compare methods. Furthermore, the best-performing method, MSG, is usually at the same distance as both the Original and Retrained model. For each method, for each initialization seed, we computed the L2 distance between the unlearned model $f_{U}$ and the retrained model $f_{R}$, as well as between the $f_{U}$ and $f_{O}$ (Figure 2).
Although having the same weight as the Retrained model would indicate that the unlearned model has unlearned $\mathcal{D}_{F}$, our evaluations show that distance to the Retrained model might not be an adequate evaluation metric for MU.
Figure 2: L2 Distance between the Unlearned ResNet18 models, the Original and Retrained models. None of the unlearned models gets close to the Retrained model’s weights; most unlearned Models are closer to the Original model than the Retrained model.
Appendix H Requirements
We ran the experiments on compute clusters with different capacities. Nonetheless, each method was tested on devices with the same specifications when recording run times: 1 NVIDIA L4 24GB GPU and 4 Intel(R) Xeon(R) CPU @ 2.20GHz.